If the two legs of a right triangle are 4 and 16 then express the hypotenuse in simplest radical form.
Answer by edjones(7569)

a^2+b^2=c^2
4^2+16^2=c^2
16+256=272
sqrt(272)=sqrt(16*17)=4sqrt(17)
.
Ed